OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2.

OpenStudy (anonymous):

Is slope-intercept form like y = mx + b ?

OpenStudy (anonymous):

y = mx + b is slope-intercept form :)

OpenStudy (anonymous):

ok, great.. they give you the slope, m, which is 3 and the y intercept, b, which is -2 So the equation is just putting in that m and b into y = mx + b y = mx + b y = 3x - 2 Do you see it now?

OpenStudy (anonymous):

yes, thank you!